Given 9.6 L of oxygen gas at STP, what is the percent yield of iron(III) oxide if 37.4 grams of iron(III) oxide is produced? Round to two decimal places and do not include a % sign in your answer.

4Fe + 3O2 ---> 2Fe2O3
1
Expert's answer
2019-03-08T05:20:18-0500

1)number of moles of oxygen = 9.6 L/ 22.4 moles/L = 0.43 moles O2

2) according to reaction stoichiometry, number of moles of iron(III) oxide produced = (0.43 moles x

2)/ 3 = 0.286 moles Fe2O3

3) mass of iron(III) oxide produced = 0.286 moles x 159.69 g/mol = 45.626 g

4) the percent yield of iron(III) oxide = (37.4 g/ 45.626 g) x 100% = 81.97%

Answer: 81.97